OpenStudy (anonymous):

1,440 is 120 percent of _________.

OpenStudy (anonymous):

use the following for all word problems is means = of means x (multiply)

OpenStudy (anonymous):

1440 = 120/100 x N ....note that all percents are put over 100

OpenStudy (anonymous):

actually the answer is 1200

OpenStudy (agreene):

another way to think about percents is to change them to decimals (this method always worked better for me). In this case: \[1400=1.2x\] solving for x: \[\frac{1400}{1/2}=x \implies x=1200 \]